These are the questions from my Math packet I can't figure out. Could someone help and explain to me how you got the answer? Thank you!

Write an equation to model the situation:  

The surface area A of a rectangular prism is the sum of the perimeter P and twice the area of the base B.

Tell whether the pairing is a function:

                  

Input

Output

-3

4

2

6

-10

8

0

5

Graph the function  with the domain x = 0, 1, 2, 3, 4, and 5.

Graph

Graph:

Solve S = 2B + Ph for B.

               

Graph the line that passes through (–5, 2) and (2, 4). Then write the equation of the line in slope-intercept form.

Write an equation of a line that is perpendicular to y = 3x – 5 and passes through (–6, 0).

Solve and graph the linear inequality: 4x + 2y > 6
Answer #1

Solve and graph the linear inequality 4x+2476 Step? divide by 2 2x+y > 3 Step 2 Solve for y y >3-2x Step3 plot in groph y=3-2

write equation of line that is perpendiculca to through (-6,0) y = 32-5 and passes Note :- two lines are perpendiculas lie me

take any point graph the line that passes through (-5,2) and (24) then write the slope - intercept ferm equation of the line

Tell me whether the paring is a function input -3 output 4 2 10 s yes its : every a function every element in input has imag

Solve for B

S = 2B + Ph

S -Ph = 2B

\frac{S-Ph}{2}= B